# How Transformers Solve Propositional Logic Problems: A Mechanistic Analysis

### v3

@@ Abstract
Large language models (LLMs) have shown amazing performance on tasks that require planning and reasoning. Motivated by this, we investigate the internal mechanisms that underpin a network’s ability to perform complex logical reasoning. We first construct a synthetic propositional logic problem that serves as a concrete test-bed for network training and evaluation. Crucially, this problem demands nontrivial planning to solve. We perform our study on two fronts. First, we pursue an understanding of precisely how a three-layer transformer, trained from scratch and attains perfect test accuracy, solves this problem. We are able to identify certain “planning” and “reasoning” mechanisms in the network that necessitate cooperation between the attention blocks to implement the desired logic. Second, we study how pretrained LLMs, namely Mistral-7B and Gemma-2-9B, solve this problem. We characterize their reasoning circuits through causal intervention experiments, providing necessity and sufficiency evidence for the circuits. We find evidence suggesting that the two models’ latent reasoning strategies are surprisingly similar, and human-like. Overall, our work systemically uncovers novel aspects of small and large transformers, and continues the study of how they plan and reason.
@@ 1 Introduction
Language models using the transformer architecture (Vaswani et al., 2017) have shown remarkable capabilities on many natural language tasks (Brown et al., 2020; Radford et al., 2019b). Trained with causal language modeling wherein the goal is next-token prediction on huge amounts of text, these models exhibit deep language understanding and generation skills. An essential milestone in the pursuit of models which can achieve a human-like artificial intelligence, is the ability to perform human-like reasoning and planning in complex unseen scenarios. While some recent works using probing analyses have shown that the activations of the deeper layers of a transformer contain rich information about certain mathematical reasoning problems (Ye et al., 2024), the question of what mechanisms inside the model enables such abilities remains unclear.
While the study of how transformers reason in general remains a daunting task, in this work, we aim to improve our mechanistic understanding of how a transformer reasons through simple propositional logic problems. For concreteness’ sake, consider the following problem:
@@
Rules: A or B implies C. D implies E. Facts: A is true. B is false. D is true.
Question: what is the truth value of C?
An answer with minimal proof is “A is true. A or B implies C; C is true.”
@@
The reasoning problem, while simple-looking on the surface, requires the model to perform several actions that are essential to more complex reasoning problems, all without chain of thought (CoT). Correctly writing down the first answer token alone takes multiple steps. It must first resolve the ambiguity of which rule is being queried relying on the question: in this case, it is “A or B implies C”. Then, it needs to rely on the premise variables A and B to resolve the ambiguity of which facts are relevant, and identify “A is true” and “B is false”. Finally, it needs to decide that “A is true” is the correct one to invoke in its answer due to the nature of disjunction.
@@
Contributions. At a high level, we make the following discoveries based on our two fronts of analysis:
  1. 1.
    3) We discover that small transformers, trained purely on the synthetic problem, utilize certain “routing embeddings” to significantly alter the information flow of the deeper layers when solving different sub-categories of the reasoning problem. We also characterize the different reasoning pathways: we find that problems querying for reasoning chains involving logical operators typically require greater involvement of all the layers in the model.
  2. 2.
@@
In particular, our analysis of Mistral-7B and Gemma-2-9B in §4 and §C.7 is, to our knowledge, the first to characterize the circuits employed by LLMs in the wild for solving a nontrivial in-context logical reasoning problem (involving distracting clauses) that requires correctly executing a multi-step reasoning chain without chain of thought. Furthermore, we make concrete progress towards general techniques of verifying the necessity and sufficiency of a pretrained LLM’s reasoning circuit.
Additionally, we define the scope of our analysis as follows. First, in the shallow transformer experiments, we focus on the variant which only has self-attention layers in addition to layer normalization, positional encoding, embedding and softmax parameters. While we could have also included MLP layers, we choose not to because the no-MLP models already achieve 100% test accuracy on the problem, and adding MLPs would unnecessarily complicate the analysis. As a second way to focus the scope of paper, in the LLM experiments, we do not seek to uncover every model component that participates in solving the reasoning problems. Instead, we focus on those which have the strongest causal influence on the model’s reasoning actions. By doing so, we can fully justify the necessity of these key components for the model’s reasoning actions, without diluting the focus with less-impactful sub-circuits (for handling low-level text processing, for instance).
1.1 Related works
Mechanistic interpretability. Our work falls in the area of mechanistic interpretability, which aims to understand the hidden mechanisms that enable language-modeling capabilities of the LLMs (Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2022; Vig et al., 2020; Feng & Steinhardt, 2024; Wu et al., 2023; Wang et al., 2023; Hanna et al., 2024; Merullo et al., 2024a; McGrath et al., 2023; Singh et al., 2024; Geiger et al., 2024; Feng et al., 2024; Marks et al., 2024; Engels et al., 2024; Merullo et al., 2024b). In particular, our work belongs to the sub-area of circuit-based mechanistic interpretability (Elhage et al., 2021; Wang et al., 2023). While the definition of a “circuit” varies across different works, in this paper, our definition is similar to the one in Wang et al. (2023): it is a collection of model components (attention heads, neurons, etc.) with the “edges” in the circuit indicating the information flow between the components in the forward pass; the “excitation” of the circuit is the input tokens.

### v4

# A Implies B: Circuit Analysis in LLMs for Propositional Logical Reasoning

@@ Abstract
Due to the size and complexity of modern large language models (LLMs), it has proven challenging to uncover the underlying mechanisms that models use to solve reasoning problems. For instance, is their reasoning for a specific problem localized to certain parts of the network? Do they break down the reasoning problem into modular components that are then executed as sequential steps as we go deeper in the model? To better understand the reasoning capability of LLMs, we study a minimal propositional logic problem that requires combining multiple facts to arrive at a solution. By studying this problem on Mistral and Gemma models, up to 27B parameters, we illuminate the core components the models use to solve such logic problems. From a mechanistic interpretability point of view, we use causal mediation analysis to uncover the pathways and components of the LLMs’ reasoning processes. Then, we offer fine-grained insights into the functions of attention heads in different layers. We not only find a sparse circuit that computes the answer, but we decompose it into sub-circuits that have four distinct and modular uses. Finally, we reveal that three distinct models – Mistral-7B, Gemma-2-9B and Gemma-2-27B – contain analogous but not identical mechanisms.
@@ 1 Introduction
LLMs can solve many tasks in a few-shot manner (Brown et al., 2020; Radford et al., 2019b). One emergent ability of these models is solving problems that require planning and reasoning. In particular, LLMs seem to have internal components that judiciously process and transform the provided information. Unfortunately, it is not clear whether these models implement step-by-step algorithms or whether they perform opaque computations to arrive at their answers.
The area of circuit analysis has emerged as a way to understand how transformers use their internal components. While the definition of a circuit varies across different works, in this paper, our definition follows Wang et al. (2023a). A circuit within a transformer is a collection of model components (attention heads, neurons, etc.) with the edges in the circuit indicating the information flow between the components in the forward pass. Circuit analysis can inform how to improve models, how to debug errors, and how to explain patterns in their behavior. Despite the importance of these goals, there has been a limited number of large-scale studies of interpreting LLM reasoning. Within mechanistic interpretability (Geva et al., 2021; Vig et al., 2020; Hou et al., 2023), existing studies either only provide partial evidence for the underlying circuits (Meng et al., 2022) or are limited to small models like a 3 layer transformer or GPT-2 sized models (Merullo et al., 2024a; Rauker et al., 2023; Wang et al., 2023a). Our goal is to go beyond this, uncovering the distinct roles of different sets of components for solving a canonical propositional reasoning problem in frontier models with as many as 27B parameters. Specifically, we aim to identify separate sub-circuits for distinct steps like retrieving information from the context, processing this information, and arriving at an answer.
1.1 A Suitable Set of Propositional Logical Reasoning Problems
Performing circuit analysis for LLMs requires a careful selection of what problem to study. If it is too simple (e.g., just copying part of the input), then we will learn little about reasoning. If it is too complex (e.g., generic math word problems), then we will have no way to systematically apply causal analysis and we will struggle to localize and delineate the model’s internal components. The problem also needs to be sufficiently natural, such that it is a common subproblem of more complex reasoning problems in the wild, such as those found in natural logic (MacCartney & Manning, 2014).
Keeping the above considerations in mind, we study the following propositional logic template, which requires reasoning over distinct parts of the input. The problem involves five boolean variables. Given two propositions as “Rules” and truth values of three variables as “Facts” we wish to infer the unknown truth value of a different variable (which we call the “query”). For concreteness, here are two possible instantiations:
@@ 1.1 A Suitable Set of Propositional Logical Reasoning Problems
Rules: $A$ or $B$ implies $C$. $D$ implies $E$. Facts: $A$ is true. $B$ is false. $D$ is true. Question: what is the truth value of $C$?     Rules: $V$ and $W$ implies $X$. $Y$ implies $Z$. Facts: $V$ is false. $W$ is true. $Y$ is false. Question: what is the truth value of $Z$?
@@
Consider the left problem, which queries the LogOp (disjunction) chain. One concise solution to the question is “A is true. A or B implies C; C is true.” This problem, while simple-looking on the surface, requires the model to perform actions that are essential to more complex reasoning problems. Even writing down the first answer token (A in “A is true”) takes multiple steps. [^1] Each word (or capital letter representing a boolean variable) gets tokenized as an individual token. The model must resolve the ambiguity of which rule is being queried. In this case, it is “A or B implies C” not “D implies E”. Then, the model needs to determine which facts are relevant and to process the information “A is true” and “B is false”. Finally, it decides that “A is true” is relevant and implies that “C is true” due to the nature of disjunction. In contrast, the problem on the right queries the linear chain, requiring the model to locate different rules ($Y$ implies $Z$) and process the facts ($Y$ is false).
Our logic problem is also inspired by some of the ones in reasoning benchmarks such as GSM8k (Cobbe et al., 2021) or ProofWriter (Tafjord et al., 2021). The issue with using existing benchmarks directly is that there is too much syntactic and semantic variation in the questions. This makes it challenging to use the tools of mechanistic interpretability, which require changing parts of the input in a systematic way. One can amplify the difficulty of our problem by increasing the number of variables or the length of the reasoning chain. However, even on this simple problem, Mistral and Gemma models only achieve 70% to 86% accuracy in writing the correct proof and determining the query value, even with few-shot prompting.
Figure 1: Left: Problem set-up. Top left shows the data model. The causal structure has two chains: one with a logical operator (LogOp, “and” or “or”) and the other with a linear causal chain. Bottom left shows how we get counterfactual signals for the model components (e.g., attention heads). We vary parts of the input, such as the query variable, truth values, and order of rules. Middle: Key steps in our interpretability approach. We use causal mediation analysis to localize components in LLMs that are used to solve the logic problem (circles represent attention heads, with layers stacked vertically). We derive evidence about the functionality of components using multiple methods. For necessity, we patch with counterfactual signals (represented by blue and red nodes). Then we check the impact of each component by looking at the output of the network (e.g., checking if it outputs very different distributions “out1” vs. “out2” after patching). For sufficiency, we replace the complement of the circuit with counterfactuals, and we verify that the output does not change too much. For fine-grained analysis, we use key-value-query interventions to identify the role each component plays. Right: New findings. Three LLMs use similar reasoning steps when solving these problems. They use four sparse and distinct sets of attention heads in a step-by-step fashion to execute rule locating, rule moving, fact processing, and decision making (depicted by the colored nodes).
1.2 Interpreting How LLMs Solve the Logic Problems
Given the above problem template, we design systematic causality-based experiments to understand how LLMs solve it. We aim to localize and delineate circuits, as a way to interpret how contemporary LLMs solve propositional logic problems (e.g., how they determine relevant parts of the input, how they move information around, how they process facts, and how they make decisions). Finding an interpretable circuit amounts to identifying a small number of components that map to natural operations and together implement an intuitive reasoning algorithm. It is important to note that, apriori, a non-interpretable outcome is also possible, where the LLM internally merges steps in reasoning or distributes the steps across all its parameters. One of our contributions is strong evidence that the mechanisms for LLMs to solve such reasoning problems are indeed modular and localized.
Figure 1 shows the main ideas of our approach, including the data model, the interpretability tools, and some findings that are shared across all three models: Gemma-2-27B, Gemma-2-9B (Gemma Team et al., 2024) and Mistral-7B-v0.1 (Jiang et al., 2023). Our main findings are:
  • All three models contain four families of attention heads with specialized roles in solving different steps in the problem: queried-rule locating heads, queried-rule mover heads, fact-processing heads, and decision heads (§3). This is surprising, as it indicates the model solves the logic problem in a sequential manner, without merging the steps. We further verify the roles of these four families by analyzing their behavior separately by patching after various types of counterfactual prompts, e.g., swapping rule locations or fact values (§3.3).
  • Intriguingly, the three LLMs all perform a type of lazy reasoning. The core circuit does not immediately pre-process the Rules and Facts in the input. Rather, it primarily activates after seeing the Question and the ultimate query variable (§3). Moreover, the LLMs tend to reuse some of their sub-circuits for different parts of the argument when possible, consistent with Merullo et al. (2024a). We verify this for Gemma-2-9B and Mistral-7B (§3.4).
  • We provide a partial analysis of Gemma-2-27B, which is somewhat limited due to computational contraints (§3.4). While all three models contain these four families components, Gemma-2-27B also possesses certain logical-operator heads that do not seem to exist or have strong causal roles in the 9B version. Gemma-2-27B’s circuit appears to be slightly more parallel than the 9B version, which is more sequential. This adds nuances to a study showing that circuits are consistent across scale by Tigges et al. (2024). Indeed, from Gemma-2-9B to Gemma-2-27B, while the core algorithm appears to stay the same, there are mechanical differences and additional functional components which likely contribute to the larger model’s higher proof accuracy.
  • Finally, we contrast pre-trained LLMs with 3-layer models trained on our task. The small models perform intermingled reasoning, linking together their attention blocks and residual streams in subtle ways. These small models have nearly perfect accuracy, but their reasoning mechanisms are less modular, combining processing facts with writing the proof. In contrast, Mistral and Gemma seem to use specialized components for different parts of the task (§5).
Our analysis is, to our knowledge, the first to characterize the circuits employed by LLMs in the wild for solving a logic problem that involves distracting clauses and requires multi-step reasoning. While this reasoning is limited compared to the abilities of today’s thinking models (Guo et al., 2025a; Arcuschin et al., 2025), we still go beyond single-hop tasks like token copying or addition from recent studies, e.g., Feucht et al. (2025); Hu et al. (2025); Yin & Steinhardt (2025). As another outcome, we have uncovered fundamental differences between small (3-layer), medium (7B/9B), and large (27B) models, indicating only some mechanistic insights generalize across scales.
2 Preliminaries and Methodology
We explain our data model and then give an overview of the methodology we use for our analysis.
@@ 2 Preliminaries and Methodology
Causal Mediation Analysis. Following prior work, we primarily rely on Causal Mediation Analysis (CMA) (Pearl, 2001) to characterize the reasoning circuits. In our setting, CMA is primarily concerned with measuring the (natural) direct and indirect effects (DE and IE) of a model component. Consider the following classic causal diagram of CMA, in Figure 2.
Figure 2: CMA causal graph.
Suppose we wish to understand whether a certain mediator $M$ plays an important role in the causal path from the input $X$ to the outcome $Y$. We decompose the “total effect” of $X$ on $Y$ into the sum of direct and indirect effects, as shown in the figure. The indirect effect (IE) measures how important a role the mediator $M$ plays in the causal path $X\to Y$. To measure it, we compute $Y$ given $X$, except that we artificially hold $M$’s output to its “corrupted” version (called an intervention), which is obtained by computing $M$ on a counterfactual (“corrupted”) version of the input. A significant change in $Y$ indicates a strong IE, which implies that $M$ is important in the causal path. On the other hand, a weak IE (so a strong DE) implies that $M$ is unimportant.
@@
Fine-grained analysis: Understanding sub-component functionality. After obtaining the full circuit, we can theorize about the role each component in the circuit plays in the reasoning chain. However, we need to verify such interpretations with empirical causal evidence. To do so, we again perform CMA with specific sub-components of the circuit held as the mediator, and we measure their IEs. For instance, if we hypothesize that an attention head uses the QUERY token to locate the rule being queried (the first reasoning step), then we perform two experiments. First, we set the mediator as the key activations of this head in the Rule section, and generate the counterfactual prompt by swapping the position of the rules, without touching anything else in the prompt. We should then observe a strong IE. Second, we generate the counterfactual prompt by flipping QUERY. Then, the query activation of this head, when held as the mediator, should have strong IE. In general, testing a hypothesis of a reasoning circuit component’s role requires careful control on where the input prompt is “corrupted”, where we need control over how the answer changes or does not change.
For the model to solve our logic problem, it has to perform multiple steps of reasoning, going from the Question to the Rules to writing down the proof. This leads to a more complicated process for systematically investigating the internal mechanisms in the LLMs. Other tools commonly used in mechanistic interpretability such as activation patching, causal tracing (Vig et al., 2020; Meng et al., 2022; Hase et al., 2024; Heimersheim & Nanda, 2024; Zhang & Nanda, 2024) are types of CMA.
3 Discovering Modular Reasoning Circuits in LLMs
To make the problem easier for the smaller models, we only examine Gemma-2-9B and Mistral-7B on the disjunction version (the LogOp is “or”); but, we study Gemma-2-27B on the full problem. For the “or” problem, Gemma-2-9B and Mistral-7B have proof accuracy around 83% and 70% respectively. Their accuracy with “and” is lower, while Gemma-2-27B has accuracy of 86% on the full problem.
We focus on Gemma-2-9B in the main text, which illustrates our analysis process. This example elucidates the common procedure we apply to the other models as well. Appendix B contains the full results for Mistral-7B and Gemma-2-27B models.
Figure 3: Gemma-2-9B’s reasoning circuit for the first answer token: The (chunks of) input tokens are on the left, which are passed into the residual stream and processed by the attention heads. We use $(\ell,h)$ to denote an attention head. When referencing multiple heads in the same layer, we write $(\ell,h_{1};h_{2};...;h_{n})$ for brevity. We illustrate the information flow manipulated by the different types of attention heads we identified to be vital to the reasoning task.
3.1 Discovering the Necessary Circuit: QUERY-based Search for Model Components
We discover the core reasoning circuit by performing QUERY-based intervention experiments. After performing attention head output patching and measuring logit differences, we locate a small set of attention heads that are central to the reasoning circuit of the LLM. The MLPs have much lower intervened logit difference (mostly $<0.1$) showing they play a limited role in the reasoning process. However, we observe MLP-0 has a higher intervened logit difference. Prior work has observed that MLP-0 acts more as a “nonlinear token embedding” than a complex high-level processing unit (Wang et al., 2023a). Hence, in the rest of this section, we focus our analysis on the attention heads.
Remark. Unless otherwise specified, we adopt a calibrated version of the logit difference (see Appendix B): the closer to 1, the stronger the indirect effect of that component is in the causal chain, and the greater a role that component plays in the reasoning circuit.
Figure 4: Patching of query and value activations in (a) and (c). We show in (b) the typical attention patterns of a representative set of the attention heads, which are important in the intervention experiments shown in (a) and (c). We make several observations from (b). Queried-rule locating head (19,11): observe that it correctly locates the queried rule that ends with Q. Queried-rule mover head (23,6): it primarily focuses on the QUERY token Q, with some small amount of attention on the tokens following it. Fact processing head (26,12): attention concentrates in the fact section. Decision head (30,9): attention focused on the correct first answer token U. In this experiment, intervening on “key” activations yields trivial scores, so we focus on “value” and “query.”.
Attention-head sub-component patching (QUERY-based patching). We now aim to understand why the attention heads identified in the last sub-section are important. For now, we continue with altering the QUERY in the prompt as our causal intervention. We intervene on the sub-components of each attention head, namely their value, key, and query, and we examine details of their attention weights. We find that there are four types of attention heads. We show the results in Figure 4.
Queried-rule locating head. Attention head (19,11)’s query activation has a large intervened logit difference according to Figure 4(a). Therefore, its query (and attention patterns) are QUERY-dependent and have strong IE. Furthermore, at the QUERY position, we find that on average, its attention weight is above 85% at the “conclusion” variable of the rule being queried. It is responsible for locating the queried rule, and storing that rule’s information at the QUERY position.222(21,7), (22,5), (23,12) exhibit similar tendencies.
Queried-rule mover head. Attention head (23,6)’s value activations have large intervened logit difference, and intriguingly, its query and key activations do not share that tendency. This already suggests that its attention pattern performs a fixed action on both the original and altered prompts, and only the value information is sensitive to QUERY. Furthermore, within the relevant context (excluding the 4 in-context examples given), (23,6) assigns around 50% attention weight to the QUERY position, and its attention weight at QUERY is about 5 times larger than the second largest one on average. Recalling the role of the shallower layers, we find evidence that (23,6) moves the QUERY and queried-rule information to the “:” position.333(20,7), (24,15) also appear to belong to this type.
Fact processing heads. Attention heads (26,12)’s query activations have large intervened logit differences. Within the relevant context, at the “:” token position, it attention weight on the sentence stating the correct fact is around 60% on average.444(24,5), (25,7), (26,0) also appear to belong to this type.
Decision head. Attention head (30,9)’s query activations have large intervened logit differences. Its attention pattern suggests that it is a “decision” head: within the relevant context, the head’s attention weight focuses on the correct Fact variable (when the model is correct). The single token position occupies about than 80% of its total attention in the relevant context on average. In other words, it locates the correct answer token.555(28,12) exhibits a similar tendency.
We delay detailed inspection and visualization of the attention statistics to Appendix B.4.2. Figure 3 illustrates the reasoning circuit we identify. We remark on two intriguing properties:
  1. nosep
  2. nosep
    There is a sparse set of attention heads that are found to be central to the reasoning circuit: the queried-rule locating heads, queried-rule mover heads, fact-processing heads, and decision heads. We discuss circuit discovery in §3.1, and verification in §3.2. This indicates that a very small fraction of the 9B parameters are taking part in solving a specific question.
3.2 Verifying Sufficiency of the Discovered Circuit
A natural question now arises: is $\mathcal{C}$ sufficient to explain the (QUERY-sensitive) reasoning actions of the LLM? We prove sufficiency by measuring the direct effect of the input prompt on the difference in the final logits, with the complement of the identified circuit treated as the mediator (defined in §2). More concretely, we run the model on the normal prompts and only allow normal information flow in every attention head in the circuit $\mathcal{C}=\{(19,11),(23,6),(26,12),(30,9),...\}$ (on or after QUERY), while freezing all the other attention heads to their counterfactual activations obtained by running on the counterfactual prompts. We expect the average circuit-intervened logit difference $\Delta^{\mathcal{C}}_{normal}$ approaching or surpassing the average logit difference of the (un-intervened) model run on the normal prompts, $\Delta_{normal}$. We confirm this hypothesis in Table 1.
$\Delta^{\mathcal{C}^{\dagger}}_{normal}/\Delta_{normal}$1.0 0.94 -0.97 -0.40 0.17 -1.11
Table 1: $\Delta^{\mathcal{C}^{\dagger}}_{normal}/\Delta_{normal}$ for Gemma-2-9B, with different choices of $\mathcal{C}^{\dagger}$. $\mathcal{C}_{null}$ denotes the empty circuit, i.e. the case where no intervention is performed. We abbreviate the attention head families, for example, $DH=$ decision heads; $\mathcal{C}-DH=$ full circuit but with the decision heads removed.
Including all 13 attention heads in $\mathcal{C}$, $\Delta^{\mathcal{C}}_{normal}$ is about 94% of $\Delta_{normal}$ on the normal samples. Removing any one of the four families of attention heads from $\mathcal{C}$ renders the direct effect almost trivial. Therefore, every head family is critical to the circuit. We present further discussions of the experimental procedure and results, and caveats of reasoning circuit verification in Appendix B.5.
3.3 Additional Evidence: More Fine-Grained Circuit Analysis
In this sub-section, we discuss example experiments of how we further verify the functionalities of the attention head families. We present the full analysis in Appendix B.8.
Queried-rule locating heads. We use the queried-rule locator heads, the “first step” in the reasoning circuit, to demonstrate how we perform finer “causal surgery” on the LLM to understand sub-components of the full circuit better.
First, based on our full-circuit CMA experiments, we have formed an interpretation of the queried-rule locator heads: they rely on the QUERY information to locate the queried rule in the Rules section. This is further corroborated by the attention statistics of these attention heads, shown in Figure 5(b)(i): these attention heads place significant amount of attention weight on the correct position of the queried rule.
To further verify the “look-up” functionality, we create counterfactual prompts by swapping the location of the rules while keeping everything else untouched in the normal prompt. For instance, we would have the following normal-counterfactual pairs:
@@ 3.3 Additional Evidence: More Fine-Grained Circuit Analysis
Clearly, this rule location swap does not cause any change to the answer (here, $C$ is true in both versions). However, the queried-rule locator heads’ functionality suggests that they heavily rely on the key activations (in the Rules section) to perform their role correctly. To provide further causal evidence for this interpretation, we should set the mediator as the key activations of the attention heads in the Rules section, and observe their indirect effects on the model’s output logits, namely the drop in the logit difference between the answer of the queried rule and that of the alternative rule.
As expected, heads with the largest IE are the queried-rule locator heads, as shown in Figure 5(a). The keys with strong indirect effects are (17,4), (19,5), (22,2) and (23,6). By noting that Gemma-2-9B uses Grouped Query Attention which results in key and value activations shared per two heads, we see that keys with strong indirect effects indeed correspond to the attention heads (17,9), (19,11), (22,5) and (23,12). In addition, we perform CMA experiments where we swap the Fact value assignments. The fact-processing heads indeed exhibit the strongest IE.
Attention statistics. We also show the weight of four attention head families in Figure 5(b)(i) to (iv). The main takeaway is correlational evidence showing that heads attend to the expected part of the input (blue bars), which varies by type of head, rather than other positions.
Figure 5: Finer-grained examination of circuit components in the Gemma-2-9B model. (a) shows the (zoomed-in) result of a finer-grained activation patching experiment, aimed at providing further causal evidence of the rule-locating heads’ functionalities. (b)(i) to (iv) show the attention statistics of the core attention head families in the circuit.
3.4 The Reasoning Circuit in Gemma-2-27B
We summarize a few interesting observations about Gemma-2-27B here (Appendix B.9 has details).
Mechanical circuit differences between 27B and 9B models. Even though the 27B model also possesses attention heads which can be roughly divided into the 4 families as in the case of the 9B model, there still are certain finer differences between their circuits: the 27B model’s circuit appears to have a greater degree of parallelism. In particular, a number of its fact-processing heads have direct effects on the model’s output: they tend to bypass the “decision” heads. On the other hand, the 9B model’s circuit is more sequential: the fact processing heads are mostly “connected” to the decision heads and exhibit weak direct effects on the model’s logits.
Logical-operator heads. As the 27B model is required to solve the full version of the problem, on top of the four generic families of attention heads, we find that there are certain “logical-operator” heads which pay particular attention to the words “and” and “or” in the rules, and exhibit strong causal influence on the model’s reasoning actions. In particular, if we generate counterfactual prompts by flipping the logical operator from “and” to “or” (or vice versa) while keeping everything else the same as the normal prompt, then we find these logical-operator heads to have strong large logit difference (hence strong indirect effects). This is further discussed in Appendix B.9.
Subcircuit reuse. Both Gemma-2-9B and Mistral-7B reuse their queried-rule locator heads and decision heads. Specifically, when the model needs to invoke a relevant rule (from the input) later to construct their output proof, it uses the same heads (but other input positions) to retrieve the location of the rule. Similarly, the decision heads may perform their function multiple times. In other words, this pair of components serve as the first and last reasoning step in both places of the model’s proof.
4 Related Work
Mechanistic Interpretability. This area explores the hidden mechanisms that enable language-modeling capabilities and other underlying phenomena from LLMs (Olsson et al., 2022; Feng & Steinhardt, 2024; Wu et al., 2023; Hanna et al., 2024; McGrath et al., 2023; Singh et al., 2024; Geiger et al., 2024; Feng et al., 2024; Marks et al., 2024; Engels et al., 2024; Merullo et al., 2024b; Todd et al., 2024). At their core, patching methods are variants of causal mediation analysis (Vig et al., 2020; Meng et al., 2022; Hase et al., 2024; Heimersheim & Nanda, 2024; Zhang & Nanda, 2024). Some work shows how models retrieve knowledge and facts (Ferrando et al., 2025; Geva et al., 2023; Sun et al., 2025; Yao et al., 2024; Wang et al., 2024b) or use anchor words (Wang et al., 2023b). Our paper is in the sub-area of circuit-based interpretability, e.g., (Elhage et al., 2021; Wang et al., 2023a). The novelty of our work is that we study larger LLMs for a problem requiring intermediate reasoning steps, which affords the identification of retrieval, processing, and decision heads.